# A mixed precision FFT with applications in MRI

### v1

@@ 2.3 Evaluation
We evaluate the numerical fidelity of the proposed mixed-precision FFT using an MRI-style workload using fully sampled complex-valued k-space data from two public datasets: fastMRI [14] (brain, 2D fast spin-echo) and SKM-TEA [15] (knee, 3D gradient recalled echo). For a typical MRI acquisition, k-space data are acquired from multiple receive array coils and individually converted to image space via FFT. A root-sum-square operation is used to get the final magnitude image. For the forward FFT this can be expressed as
and the round-trip FFT per coil is expressed as,
where is a 2-D single coil element in the frequency domain, is a 2-D coil element in the image domain and and are our proposed mixed-precision forward and inverse FFT implementation respectively. We compare our implementation against the built-in numpy FP32 FFT reference and our FP16 FFT positive control. The peak-signal-to-noise ratio (PSNR), structural similarity index (SSIM) and normalized mean squared error (NMSE) were evaluated across 10 RSS images from both datasets. Figure 1A shows the overall experimental setup.

### v2

@@ 2.3 Evaluation
We evaluate the numerical fidelity of the proposed mixed-precision FFT using an MRI-style workload using fully sampled complex-valued k-space data from two public datasets: fastMRI [19] (brain, 2D fast spin-echo) and SKM-TEA [2] (knee, 3D gradient recalled echo). For a typical MRI acquisition, k-space data are acquired from multiple receive array coils and individually converted to image space via FFT. A root-sum-square operation is used to get the final magnitude image. The forward and round trip FFT per coil can be expressed as
where is a 2-D single coil element in the frequency domain, is a 2-D coil element in the image domain and and are our proposed mixed-precision forward and inverse FFT implementation respectively. We compare our implementation against the built-in numpy FP32 FFT reference and our FP16 FFT positive control. Standard computational imaging/computer vision metrics peak-signal-to-noise ratio (PSNR), structural similarity index (SSIM [18]) and normalized mean squared error (NMSE) were evaluated across 10 RSS images from both datasets. Figure 1A shows the overall experimental setup.